OpenStudy (osprey):

4-4 Our galaxy is about 10^5 light years across, and the most energetic particles known have an energy of about 10^19eV. How long would it take a proton with this energy to traverse the galaxy as measured in the rest frame of The galaxy ? (10^5 yrs) The particle ? (5 mins)

OpenStudy (osprey):

It's the 10^19eV that SEEMS to be causing me problems here. The numbers I get seem to go into a lot of decimal places just shy of c. and my calculator and brain explode ...

OpenStudy (irishboy123):

Using rest energy of a proton \( m_p = 938.272 MeV/c^2\), we can see \(\gamma = \dfrac{E}{E_0} = \dfrac{10^{19}}{938.272 \cdot 10^6} \approx 1.07 \cdot 10^{10}\) With events: - \(E_1\): proton starts journey, - \(E_2\) proton finishes journey, ...... we have space-time co-ordinates: S: rest frame of Universe \(E_1: (0,0)\) \(E_2: (x,t)\) S': rest frame of proton \(E'_1: (0,0)\) \(E'_2: (0,t')\) In frame S, it is simply: \(t = \dfrac{ x}{v}\) And given \(\beta = \sqrt { 1- \dfrac{1}{\gamma^2} } \approx 1\) then \(t = \dfrac{ 10^5 c }{c} = 10^5\) In frame S', firrstly, the Lorentx transforms have simplified because \(x' = 0\) to these: \(x = \gamma v t'\) and \(t = \gamma t'\) from the second \(t' = \dfrac{t}{\gamma } = \dfrac{10^5 * 365 * 24 * 60}{1.07 * 10^{10} } \approx 4.96 \) min in terms of calculator......i used the net :-)) https://www.wolframalpha.com/input/?i=(10%5E5+*+365+*+24+*+60)%2F(1.06+*+10%5E(10)+) https://www.wolframalpha.com/input/?i=((10%5E19)%2F(938.272+*+10%5E6)+)